OpenStudy (anonymous):

Solve by factoring: m^2+8m+7=0 Which answer is correct....(A. 8, 7),(B. -7, 1),(C. -7, -1), or (D. 7, 1)

OpenStudy (anikay):

Okay so we're factoring. do you know the easiest way to factor?

OpenStudy (anikay):

by completing the square.

OpenStudy (anonymous):

The square?

OpenStudy (anikay):

mmk, let me explain it's impossible to factor the expression you gave me, as it is.

OpenStudy (anikay):

so we mess with the equation a little and "complete to square" to enable factoring.

OpenStudy (anikay):

so what I'm saying is \[m^2+8m+7=0\] cannot be factored we must instead complete the square \[m^2+8m+7+9=0+9\]

OpenStudy (anonymous):

Ah

OpenStudy (anikay):

so give it a try. see if you can do it from there.

OpenStudy (anonymous):

multiple choice supposes that one of the answers is correct to start with. To save time on a test you can simply dbl chk the options to see what fits. Otherwise, anikay's process is fine :)